OpenStudy (anonymous):

Solve 6x – y = 1 for y. Is the answer: y= -6x+1??? Im not sure.. thank u

OpenStudy (anonymous):

6x-y=1 -y=-6x+1 (subtract 6 from both sides) y=6x-1 (divide both sides by -1 to get y alone)

OpenStudy (anonymous):

thanks